OpenStudy (anonymous):

How do you get this to equal Zero?? -3(x+9)^=-63

myininaya (myininaya):

the expression looks a bit weird are you sure this is how it looks?

OpenStudy (anonymous):

yes... -_____-

OpenStudy (anonymous):

It's squared... the ^ thinggyy? idk. im dumb

myininaya (myininaya):

no you aren't so we have -3(x+9)^2=-63

myininaya (myininaya):

i will write that in pretty latex

myininaya (myininaya):

\[-3(x+9)^2=-63\]

myininaya (myininaya):

First of all if you write this as something=0 then you are making the problem more hard

OpenStudy (anonymous):

yah, that

myininaya (myininaya):

All you have to do is first divide -3 on both sides like so \[(x+9)^2=\frac{63}{3}\]

myininaya (myininaya):

\[(x+9)^2=21\]

myininaya (myininaya):

now to get rid of the square you need to take square root of both sides

myininaya (myininaya):

\[\sqrt{(x+9)^2}=\sqrt{21}\]

myininaya (myininaya):

But this is the same thing as saying \[|x+9|=21\]

myininaya (myininaya):

\[=> x+9=\pm 21\]

myininaya (myininaya):

Finally subtract 9 on both sides to find x

myininaya (myininaya):

\[x=\pm 21-9\]

myininaya (myininaya):

this means you have two solutions \[x=21-9 \text{ or } x=-21-9\]

OpenStudy (anonymous):

in the begining why wouldnt you distribute the -3 ?

OpenStudy (anonymous):

i need the problem where i can put it in the quadratic formula

myininaya (myininaya):

\[-3(x^2+18x+81)=-63\] divide both sides by -3 \[(x^2+18x+81)=21\] \[x^2+18x+81=21\] now subtract 21 on both sides \[x^2+18x+60=0\]